OpenStudy (nbarrera):

Solve. z + 17 = 9 –9 –8 8 9

OpenStudy (anonymous):

whats 17-9?

OpenStudy (nbarrera):

thx

OpenStudy (shadowlegendx):

z + 17 = 9 Subtract 17 from both sides z = -8

OpenStudy (confluxepic):

The first step is to subtract.

OpenStudy (anonymous):

-8

OpenStudy (confluxepic):

z+17= 9 -17 -17 z=9-17 z=?

OpenStudy (confluxepic):

@LeeEtchison You are supposed to guide them to the answer not tell them the answer. Somebody just did that and a mod deleted their response and probably suspended them.

OpenStudy (anonymous):

Oh..sorry, I'm new here..